OpenStudy (drewbooks):

Need help with simplifying this: -3(10x + 4y) + 6(6x - 2y)

OpenStudy (drewbooks):

These are the possible answers: A) -18x^2y^2 B) 6x - 24y C) 6x^2 - 24y^2 D) 6x + 2y

OpenStudy (anonymous):

6x -24y, B.

OpenStudy (drewbooks):

Would you mind explaining how you got to that?

OpenStudy (anonymous):

Let's simplify step-by-step. −3(10x+4y)+6(6x−2y) Distribute: =(−3)(10x)+(−3)(4y)+(6)(6x)+(6)(−2y) =−30x+−12y+36x+−12y Combine Like Terms: =−30x+−12y+36x+−12y =(−30x+36x)+(−12y+−12y) =6x+−24y

OpenStudy (anonymous):

Answer: =6x−24y, B.

OpenStudy (drewbooks):

Cheers. :-)